OpenStudy (anonymous):

x^2 -4x+85=0 The answer choices are 7+3i, 7-3i 3+7i, 3-7i 2+19i, 2-19i 2+9i, 2-9i

OpenStudy (blank ):

I got x = 2+9i, 2-9i The answer is D. 2+9i, 2-9i

OpenStudy (blank ):

I could show you the steps if you want to see how I got that.

OpenStudy (solomonzelman):

\[x^2 -4x+85=0\]\[x^2 -4x=-85\]\[x^2 -4x+4=-85+24\]\[x^2 -4x+4=-81\]\[(x-2)^2=-81\]\[x-2=± \sqrt{-81}\]\[x-2=±9i\]\[x=2±9i\]

OpenStudy (blank ):

Yep ^